OpenStudy (anonymous):

Solve the system algebraically. Check your work. 5x + 2y = 10 3x + 2y = 6

OpenStudy (arindameducationusc):

i m here

OpenStudy (anonymous):

x=2 y=0

OpenStudy (usukidoll):

the y's seems to be the same so we can distribute a negative sign all over one equation and use elimination method to solve for x.

OpenStudy (usukidoll):

@saseal please don't give direct answers as it's a violation of OpenStudy's Code of Conduct.

OpenStudy (anonymous):

oh ok..

OpenStudy (anonymous):

Try using matrix method. It works everytime.

OpenStudy (arindameducationusc):

Check this

OpenStudy (usukidoll):

for some reason matrix method is taught way later :(

OpenStudy (anonymous):

Yea, they are saving the best for the last lol.

OpenStudy (arindameducationusc):

x John can find out now, what say guys?

OpenStudy (usukidoll):

?

OpenStudy (arindameducationusc):

Y is in the solution I gave, now x he can easily find out by putting in the equation

OpenStudy (usukidoll):

oh XD

OpenStudy (anonymous):

\[A=\left[\begin{matrix}5 & 2 \\ 3 & 2\end{matrix}\right]\] \[B = \left(\begin{matrix}x \\ y\end{matrix}\right)\] \[C = \left(\begin{matrix}10 \\ 6\end{matrix}\right)\] \[A ^{-1} = \left[\begin{matrix}0.5 & -0.5 \\ -0.75 & 1.25\end{matrix}\right]\] \[A ^{-1}C = B = \left[\begin{matrix}0.5 & -0.5 \\ -0.75 & 1.25\end{matrix}\right] \left(\begin{matrix}10 \\ 6\end{matrix}\right) = Find out yourself :)\]